In a certain plant species, inheritance of fruit size is controlled by four genes, each of which has two alleles that show incomplete dominance and each of which contributes equally to an individual’s phenotype. Plants with the largest fruits have the genotype AABBCCDD, while those with the smallest fruits have the genotype aabbccdd.

a) How many possible genotypes for fruit size are there in this species?
b) Based on genotype alone (ignoring environmental factors), how many possible phenotypes for fruit size are there in this species?
c) In a cross of an individual with the largest possible fruits by an individual with the smallest possible fruits, how many phenotypes will be observed in the F1 generation?
d) What will be the phenotype of an individual heterozygous for all four genes? What percentage of the offspring from a cross of two such individuals will have the largest possible fruits? What percentage of the offspring will have the smallest possible fruits?
e) What will be the phenotype of an individual with the genotype AABBccdd? What will be the phenotype of an individual with the genotype aabbCCDD? What percentage of the offspring from a cross of these two individuals will have the largest possible fruits? What percentage of the offspring will have the smallest possible fruits?
Answer #1

Answer :

(a) Here we have four genes (A, B, C, D) , thus n=4.

Number of possible genotype = 3^n

=3^(4)

=81 genotypes.

(b) Possible number of phenotypes in case of incomplete dominance: 2n + 1

= (2×4) +1

= 9 phenotypes.

(c) If we cross an individual with largest fruit size (AABBCCDD) with an individual with smallest fruit size (aabbccdd) then all offspring will be AaBbCcDd and a single phenotype will be observed in the F1 generation.

(d) Phenotype of an individual heterozygous for all four genes : AaBbCcDd.

Upon crossing such two individual (AaBbCcDd×AaBbCcDd):

  • The percentage for individuals with largest fruit size -

= p(AAABBCCDD) × 100%

= p(AA) × p(BB) × p(CC) × p(DD) × 100

= (1/4) ×(1/4) ×(1/4) ×(1/4) ×100

= 0.39%

  • The percentage of individuals with smallest possible fruits (aabbccdd) will be same as that of AABBCCDD, 0.39%.

(e) Phenotype of an individual with either AABBccdd or aabbCCDD will be a hybrid one, with intermediate size (half the size of largest possible sized fruit ).

Upon crossing AABBccdd × aabbCCDD, we will have AaBbCcDd individuals in the offspring and all will have intermediate size.

Thus, percentage of largest or smallest sized fruits will be zero.